# Rankability-enhanced Revenue Uplift Modeling Framework for Online Marketing

### v1

@@ 4.3. Response Ranking Learning
From the above analysis in Sec. 4.2, we obtain two tighter uplift modeling error bounds from the ranking perspective. In this section, we transform them into the within-group response ranking loss and cross-group ranking loss between the predicted responses and the real responses, respectively, to enhance the uplift model’s rankability.
@@ 4.3.1. Within-group Response Ranking Loss
In Eq. 10, for the individual pair $(i,j)$, if the relative order between predicted responses differs from that between true responses, we punish the uplift model with the above alignment loss. Adding the within-group response ranking loss in the treatment group to that in the control group, we can obtain the overall $\mathcal{L}_{wr-rank}$:
@@ 4.3.2. Cross-group Response Ranking Loss
Similarly, we punish the uplift model when the predicted responses of individual pair $(i,j)$ cannot meet the cross-group response-related order criteria. Then, adding the treatment-control response ranking loss to the control-treatment response ranking loss, we obtain:
@@ 4.4. Uplift Ranking Learning
In this section, we directly dive into the uplift ranking problem and explicitly optimize the uplift model’s rankability. It can be easily noted that this is inherently a learning-to-rank problem. The previous solutions in this area mainly include three types of approaches: pointwise, pairwise, and listwise. The former two kinds consider the ranking problem from the perspective of a single individual or an individual pair, thus can hardly take all individuals into consideration at a glance. Besides, the pairwise approaches are mainly classification-based and thus have great difficulty distinguishing the inner-pair uplift distance among different pairs. Therefore, they are more suitable for binary responses, i.e., conversion uplift problems. If deploying them in the revenue uplift scenario where the response is continuous, the performance can hardly be satisfying. The listwise approach, however, can directly minimize the overall ranking error of all individuals in the population and discern the difference in continuous revenue uplift values corresponding to different individuals.
@@ 4.4.2. Listwise Uplift Ranking Loss
With the above definition, given the list of true uplift score $\tau$ and the list of predicted uplift score $\hat{\tau}$ by the model, we can measure the distance between them as the overall ranking error. Here, we utilize the Cross Entropy as the measurement and define the listwise uplift ranking loss as follows:
@@
Considering that $\sum_{\mathit{X}}\tau(\mathit{X})$ is actually an unknown constant for each instance of variable $\mathit{X}$, we remove it from the equation to facilitate the implementation.
@@
Under the setting of RCT experiments, the assumption that $\mathit{X},\mathit{X}^{t},\mathit{X}^{c}$ follow the same distribution holds. Thus, we can obtain:
@@ 4.5. Overall Training Objective
where $\lambda$ is the coefficient of the L2 regularization term.
@@ 5. Experiments
  • RQ4: How does our method perform in the online deployment scenario compared with other online service strategies?
5.1. Experiment Setup
5.1.1. Datasets
5.1.2. Baselines
@@ 5.1.3. Evaluation Metrics
In our experiments, we mainly adopt the following four metrics to evaluate the uplift ranking ability of different methods: Area Under the uplift curve (AUUC), Area Under the QINI Curve (AUQC), Kendall Rank Correlation Coefficient (KRCC), and LIFT@h ($h$ is set as 30). Following the previous work (Devriendt et al., 2020), we take the “jointly, absolute” version of uplift/Qini curves and further formulate the corresponding AUUC and AUQC metrics. For easier computation and fairer comparison, we utilize the normalized AUUC and AUQC approximated over 100 buckets. The detailed evaluation metric introduction is provided in Appendix C.
We provide the implementation details in Appendix D and source codes in https://github.com/BokwaiHo/revenue_uplift.
@@
5.2. Overall Performance (RQ1)
@@ 5.3. Ablation Study (RQ2)
5.4. Hyperparameter Sensitivity Analysis (RQ3)
@@
5.5. Online Deployment (RQ4)
5.5.1. Online Deployment Scenario Description
@@
5.5.2. Online Experimental Results Analysis
From the results shown in Table 3, we can observe that our RERUM achieves consistent improvement over the state-of-the-art (SOTA) online base model across three campaigns. The average improvement reaches $20.61\%$ on the sales revenue LIFT@2 metric. Besides, the RERUM brings 430 million dollar assets under management (AUM) gain each month. Such empirical results strongly demonstrate the effectiveness of our RERUM in real-world applications.
@@ 5.5.2. Online Experimental Results Analysis
6. Conclusions
Revenue uplift modeling has been recognized as of great importance for online marketing. Especially, the uplift ranking among the whole population is the core to identifying the sensitive individuals to interventions. In this paper, we have three contributions to address this problem: 1) modeling the response regression with a ZILN loss and a customized network framework that can be adapted with different base uplift models; 2) analyzing the uplift ranking errors and using two tighter response ranking losses to augment the vanilla regression term; 3) directly modeling the listwise uplift ranking among the whole population. The extensive experiments demonstrate the effectiveness of our proposed method.
@@ Appendix C Descriptions to Evaluation Metrics
Appendix D Implementation Details
We conduct each experiment with five different random seeds and take the average performance as the final result. We use the Adam optimizer with the initial learning rate as 0.001. The embedding size for each feature is set as 10 following the conventional design in industrial applications. We implement our method with PyTorch and run it on an NVIDIA Tesla V100 GPU with 32GB memory.
Table 5. Comparison among different versions of RERUM with various response regression modules. The combination of Huber loss and exponential distribution, Huber loss and log-cauthy distribution are abbreviated as Huber(Exo) and Huber(LC), respectively.
Methods Hillstrom-Men Hillstrom-Women Product
AUUC AUQC KRCC LIFT@30 AUUC AUQC KRCC LIFT@30 AUUC AUQC KRCC LIFT@30
RERUM (TAR, Huber(Exo)) 0.6078 0.6146 0.1254 0.9642 0.6038 0.6065 0.1337 0.7910 0.6607 0.6788 0.3117 0.9451
RERUM (TAR, Huber(LC)) 0.6112 0.6181 0.1376 1.0531 0.6176 0.6189 0.1509 0.8128 0.6712 0.6874 0.3138 0.9642
RERUM (TAR, ZILN) 0.6299 0.6338 0.1617 1.3477 0.6360 0.6334 0.1806 0.8755 0.6863 0.7086 0.3269 1.0247
RERUM (CFRwass, Huber(Exo)) 0.6165 0.6151 0.1194 1.1238 0.6235 0.6283 0.1217 0.7623 0.7381 0.7367 0.3357 1.0254
RERUM (CFRwass, Huber(LC)) 0.6207 0.6233 0.1208 1.2074 0.6304 0.6354 0.1312 0.7718 0.7405 0.7424 0.3419 1.0456
RERUM (CFRwass, ZILN) 0.6474 0.6497 0.1382 1.4602 0.6568 0.6559 0.1601 0.8261 0.7563 0.7679 0.3625 1.1302
RERUM (CFRmmd, Huber(Exo)) 0.6072 0.6097 0.1158 1.1239 0.6108 0.6145 0.1125 0.7413 0.7260 7309 0.3465 0.9623
RERUM (CFRmmd, Huber(LC)) 0.6113 0.6161 0.1372 1.1626 0.6172 0.6208 0.1407 0.7762 0.7322 0.7438 0.3517 1.0724
RERUM (CFRmmd, ZILN) 0.6242 0.6281 0.1581 1.3015 0.6374 0.6350 0.1692 0.8252 0.7554 0.7647 0.3588 1.2045
RERUM (DragonNet, Huber(Exo)) 0.6431 0.6473 0.1145 1.4719 0.6257 0.6230 0.1496 0.8641 0.6758 0.6801 0.3502 0.9928
RERUM (DragonNet, Huber(LC)) 0.6525 0.6544 0.1232 1.4950 0.6324 0.6299 0.1527 0.9028 0.6792 0.6905 0.3567 1.0233
RERUM (DragonNet, ZILN) 0.6721 0.6753 0.1434 1.5845 0.6580 0.6566 0.1664 0.9401 0.7176 0.7359 0.3653 1.1016
Appendix E Empirical Analysis to Response Regression Loss Selection
In the method design, we adopt the zero-inflated lognormal (ZILN) loss for response regression, which is based on two observations: 1) zero values occupy the main frequency of the revenue response distribution; 2) the remaining positive responses approximately follow a lognormal distribution. Though its effectiveness has been comprehensively demonstrated in our Sec. 5, this customized design may be argued to be ad hoc to some extent.
Therefore, to further verify the optimal suitability of our ZILN regression module, we also investigate other potential schemes ro regress revenue responses. Especially, we replace the ZILN module in the overall RERUM framework with the combination of Huber loss (Huber, 1992) and exponential/log-cauthy distributions. The Huber loss exhibits the advantage of not being outliers sensitive, and the zero values in our revenue responses can be regarded as outliers. The exponential or log-cauthy distribution may also be a kind of reasonable distribution to model the positive revenue responses. The related empirical results comparison are provided in the Table 5.
From the results in the table, we can find that the version of our framework with ZILN can obviously ourperform the versions with other regression losses, which demonstrates the effectiveness and rationality of utilizing ZILN as revenue response regression loss.
Methods
Hillstrom-
Men
Hillstrom-
Women
Product
TAR 0.2142 0.2087 0.4429
TAR+ZILN 0.0586 0.0514 0.1378
TAR+ZILN+UR 0.0614 0.0575 0.1476
TAR+ZILN+RR 0.0608 0.0550 0.1423
RERUM (TAR) 0.0667 0.0634 0.1535
CFRwass 0.2357 0.2024 0.3966
CFRwass+ZILN 0.0651 0.0488 0.1102
CFRwass+ZILN+UR 0.0685 0.0524 0.1197
CFRwass+ZILN+RR 0.0677 0.0506 0.1174
RERUM (CFRwass) 0.0741 0.0589 0.1265
CFRmmd 0.2456 0.2120 0.4251
CFRmmd+ZILN 0.0712 0.0634 0.1206
CFRmmd+ZILN+UR 0.0785 0.0697 0.1385
CFRmmd+ZILN+RR 0.0764 0.0675 0.1310
RERUM (CFRmmd) 0.0823 0.0731 0.1436
DragonNet 0.1955 0.2037 0.3754
DragonNet+ZILN 0.0483 0.0528 0.0986
DragonNet+ZILN+UR 0.0545 0.0617 0.1028
DragonNet+ZILN+RR 0.0516 0.0574 0.1001
RERUM (DragonNet) 0.0628 0.0654 0.1109
@@ Methods
Appendix F Discussions on Impacts of Different Modules to Response Regression
As emphasized in the Sec. 1, the focus of this work is enhancing the uplift model’s rankability and all three proposed modules are surrounding how to boost the model’s capability to identify the most susceptible individuals to the treatment. Even utilizing ZILN instead of the conventional MSE to conduct the regression is also for improving ranking performance via accurate predicting the revenue response. Even though, investigating how such modules (especially UR and RR modules which are not specially designed for response regression) fare on the underlying response regression task can help provide the more comprehensive understanding to our methods. Therefore, we compare the performance of X (X indicates the base model), X+ZILN, X+ZILN+UR, RERUM(X), on the commonly used regression metric Mean Absolute Percentage Error (MAPE) of the response prediction (lower is better). Note that all base models X utilize the conventional MSE as the response regression loss according to their original papers. The detailed experimental results are provided in Table 6.
From the Table 6, we can find that X with ZILN as the regression loss can significantly outperform the X with the MSE as the regression loss on the MAPE metric. This demonstrates that the ZILN loss can effectively help predict the revenue responses more accurately, thus finally benefiting the uplift ranking. Besides, the introduction of UR and RR will influence the performance on MAPE metric a little bit. Because their main goal is to increase the uplift ranking accuray instead of the response regression. As long as they can facilitate the uplift ranking performance (shown in the Table 1 of the paper), a small degree of decrease in response regression is still tolerable. Even though, the regression performance (MAPE) of X+ZILN+UR, X+ZILN+RR, RERUM(X) is still better than X with MSE as the regression loss.
Appendix G Supplementary Ablation Study
In addition to the ablation study conducted in Sec. 5.3, we also adjust the module stacking order to further highlight the effect of each module to the revenue uplift modeling. In detail, we incrementally add the ZILN regression module (ZILN), uplift ranking learning module (UR), and response ranking learning module (RR) to the vanilla base model X. The corresponding experimental results are illustrated in Table 7. First, it is evident that X+ZILN can achieve better performance than vinalla X under almost all ranking-related metrics and datasets, which implies that ZILN module can indeed improve the rankability of uplift models. This is actually consistent with our discussion in Sec. 4.1. Second, we can also notice that integrating each module into the overall framework can respectively bring about improvement in ranking performance, which further verifies their effectiveness.
@@ Appendix G Supplementary Ablation Study
Appendix H Pseudocode
To better illustrate the process flow of our proposed RERUM framework, we provide the pseudocode as follows in Algorithm 1.
1 Process:;
3 while Not Converged do
8             Sample $S$ individuals from $\mathcal{B}^{t}$ and $\mathcal{B}^{c}$ separately and compute $\mathcal{L}^{tc}_{cr-rank}$ and $\mathcal{L}^{ct}_{cr-rank}$ with Eqn. 11 ;
12
13       end for
14
15 end while
Algorithm 1 RERUM

### v2

@@ 4.3. Response Ranking Learning
From the above analysis in Sec. 4.2, we obtain two tighter uplift modeling error bounds from the ranking perspective. Thus, we can transform them into the within-group response ranking loss and cross-group ranking loss between the predicted and the real responses, respectively, to enhance the uplift model’s rankability.
@@ 4.3.1. Within-group Response Ranking Loss
In Eq. 10, for the individual pair $(i,j)$, if the relative order between predicted responses differs from that between true responses, we punish the uplift model with the above alignment loss. Adding the within-group response ranking loss in the treatment group $\mathcal{L}^{tt}_{wr-rank}$ to the within-group response ranking loss in the control group $\mathcal{L}^{cc}_{wr-rank}$, we can obtain the overall $\mathcal{L}_{wr-rank}$:
@@ 4.3.2. Cross-group Response Ranking Loss
Similarly, we punish the uplift model when the predicted responses of individual pair $(i,j)$ cannot meet the cross-group response-related order criteria. Then, adding the treatment-control response ranking loss $\mathcal{L}^{tc}_{cr-rank}$ to the control-treatment response ranking loss $\mathcal{L}^{ct}_{cr-rank}$, we obtain:
@@ 4.4. Uplift Ranking Learning
In this section, we directly dive into the uplift ranking problem and explicitly optimize the uplift model’s rankability. It can be easily noted that this is inherently a learning-to-rank problem. The previous solutions mainly include three types: pointwise, pairwise, and listwise. The former two kinds consider the ranking problem from the perspective of a single individual or an individual pair, thus can hardly take all individuals into consideration at a glance. Besides, the pairwise approaches are mainly classification-based and thus have great difficulty distinguishing the inner-pair uplift distance among different pairs. Therefore, they are more suitable for binary responses, i.e., conversion uplift problems. If they are deployed in the revenue uplift scenario where the response is continuous, the performance can hardly be satisfying. The listwise approach, however, can directly minimize the overall ranking error of all individuals in the population and discern the difference in continuous revenue uplift values corresponding to different individuals.
@@ 4.4.2. Listwise Uplift Ranking Loss
With the above definition, given the list of true uplift score $\tau$ and the list of predicted uplift score $\hat{\tau}$ by the model, we can measure the distance between them as the overall ranking error. Here, we utilize the Cross-Entropy as the measurement and define the listwise uplift ranking loss as follows:
@@
Considering that $\sum_{\mathit{X}}\tau(\mathit{X})$ is an unknown constant for each instance of variable $\mathit{X}$, we remove it from the equation to facilitate the implementation.
@@
Under the setting of RCT experiments, the assumption that $\mathit{X}$, $\mathit{X}^{t}$, and $\mathit{X}^{c}$ follow the same distribution holds. Thus, we can obtain:
@@ 4.5. Overall Training Objective
where $\lambda$ is the coefficient of the L2 regularization term. To better illustrate the process flow of our proposed RERUM framework, we provide the pseudocode in Algorithm 1.
1 Process:;
3 while Not Converged do
8             Sample $S$ individuals from $\mathcal{B}^{t}$ and $\mathcal{B}^{c}$ separately and compute $\mathcal{L}^{tc}_{cr-rank}$ and $\mathcal{L}^{ct}_{cr-rank}$ with Eqn. 12 ;
12
13       end for
14
15 end while
Algorithm 1 RERUM
@@ 5. Experiments
  • RQ4: How do our proposed modules influence the response regression performance?
  • RQ5: How does our method perform in the online deployment scenario compared with other online service strategies?
@@ 5.1.3. Evaluation Metrics
In our experiments, we adopt the following four metrics to evaluate the uplift ranking ability of different methods: Area Under the uplift curve (AUUC), Area Under the QINI Curve (AUQC), Kendall Rank Correlation Coefficient (KRCC), and LIFT@h ($h$ is set as 30). Following the previous work (Devriendt et al., 2020), we take the “jointly, absolute” version of uplift/Qini curves and further formulate the corresponding AUUC and AUQC metrics. For easier computation and fairer comparison, we utilize the normalized AUUC and AUQC approximated over 100 buckets. The detailed evaluation metric introduction is provided in Appendix C.
5.1.4. Implementation Details
We conduct each experiment with five different random seeds and take the average performance as the final result. We use the Adam optimizer with the initial learning rate as 0.001. The embedding size for each feature is set as 10 following the conventional design in industrial applications. We implement our method with PyTorch and run it on an NVIDIA Tesla V100 GPU with 32GB memory. Besides, we provide source codes in https://github.com/BokwaiHo/revenue_uplift.
@@ 5.1.4. Implementation Details
5.2. Results and Analysis
5.2.1. Overall Performance (RQ1)
@@ 5.2.2. Ablation Study (RQ2)
5.2.3. Hyperparameter Sensitivity Analysis (RQ3)
@@
5.2.4. Response Regression Performance Analysis (RQ4)
Methods
Hillstrom-
Men
Hillstrom-
Women
Product
TAR 0.2142 0.2087 0.4429
TAR+ZILN 0.0586 0.0514 0.1378
TAR+ZILN+UR 0.0614 0.0575 0.1476
TAR+ZILN+RR 0.0608 0.0550 0.1423
RERUM (TAR) 0.0667 0.0634 0.1535
CFRwass 0.2357 0.2024 0.3966
CFRwass+ZILN 0.0651 0.0488 0.1102
CFRwass+ZILN+UR 0.0685 0.0524 0.1197
CFRwass+ZILN+RR 0.0677 0.0506 0.1174
RERUM (CFRwass) 0.0741 0.0589 0.1265
CFRmmd 0.2456 0.2120 0.4251
CFRmmd+ZILN 0.0712 0.0634 0.1206
CFRmmd+ZILN+UR 0.0785 0.0697 0.1385
CFRmmd+ZILN+RR 0.0764 0.0675 0.1310
RERUM (CFRmmd) 0.0823 0.0731 0.1436
DragonNet 0.1955 0.2037 0.3754
DragonNet+ZILN 0.0483 0.0528 0.0986
DragonNet+ZILN+UR 0.0545 0.0617 0.1028
DragonNet+ZILN+RR 0.0516 0.0574 0.1001
RERUM (DragonNet) 0.0628 0.0654 0.1109
As emphasized in the Sec. 1, the focus of this work is enhancing the uplift model’s rankability and all three proposed modules are surrounding how to boost the model’s capability to identify the most susceptible individuals to the treatment. Even utilizing ZILN instead of the conventional MSE to conduct the regression is also for improving ranking performance via accurate predicting the revenue response. Even though, investigating how such modules (especially UR and RR modules which are not specially designed for response regression) fare on the underlying response regression task can help provide the more comprehensive understanding to our methods. Therefore, we compare the performance of X (X indicates the base model), X+ZILN, X+ZILN+UR, RERUM(X), on the commonly used regression metric Mean Absolute Percentage Error (MAPE) of the response prediction (lower is better). Note that all base models X utilize the conventional MSE as the response regression loss according to their original papers. The detailed experimental results are provided in Table 3. From the Table 3, we can find that X with ZILN as the regression loss can significantly outperform the X with the MSE as the regression loss on the MAPE metric. This demonstrates that the ZILN loss can effectively help predict the revenue responses more accurately, thus finally benefiting the uplift ranking. Besides, the introduction of UR and RR will influence the performance on MAPE metric a little bit. Because their main goal is to increase the uplift ranking accuray instead of the response regression. As long as they can facilitate the uplift ranking performance (shown in the Table 1 of the paper), a small degree of decrease in response regression is still tolerable. Even though, the regression performance (MAPE) of X+ZILN+UR, X+ZILN+RR, RERUM(X) is still better than X with MSE as the regression loss.
5.2.5. Online Deployment (RQ5)
@@
Online Experimental Results Analysis. From the results shown in Table 4, we can observe that our RERUM achieves consistent improvement over the state-of-the-art (SOTA) online base model across three campaigns. The average improvement reaches $20.61\%$ on the sales revenue LIFT@2 metric. Besides, the RERUM brings 430 million dollar assets under management (AUM) gain each month. Such empirical results strongly demonstrate the effectiveness of our RERUM in real-world applications.
@@
6. Conclusions and Future Work
Revenue uplift modeling has been recognized as of great importance for online marketing. Especially, the uplift ranking within the whole population is the core to identifying the sensitive individuals to interventions. However, previous methods suffer from the low revenue response prediction precision and the neglection to the uplift ranking accuracy among individuals. In this paper, we have three contributions to address this problem: 1) modeling the response regression with a ZILN loss and a customized network framework that can be adapted with different base uplift models; 2) analyzing the uplift ranking errors and using two tighter response ranking losses to augment the vanilla regression term; 3) directly modeling the listwise uplift ranking among the whole population. The extensive experiments on both offline datasets and one of the world’s largest online fintech marketing platforms demonstrate the effectiveness of our proposed method. In the future, we prepare to explore the revenue uplift modeling with multiple treatments and its application in cross-domain scenarios.
Acknowledgements.
This work was supported by the Start-up Grant (No. 9610564), the Strategic Research Grant (No. 7005847) of the City University of Hong Kong, and the Early Career Scheme (No. CityU 21219323) of the University Grants Committee (UGC).
@@ Appendix C Descriptions to Evaluation Metrics
Appendix D Supplementary Ablation Study
In addition to the ablation study conducted in Sec. 5.2.2, we also adjust the module stacking order to further highlight the effect of each module to the revenue uplift modeling. In detail, we incrementally add the ZILN regression module (ZILN), uplift ranking learning module (UR), and response ranking learning module (RR) to the vanilla base model X. The corresponding experimental results are illustrated in Table 6. First, it is evident that X+ZILN can achieve better performance than vinalla X under almost all ranking-related metrics and datasets, which implies that ZILN module can indeed improve the rankability of uplift models. This is actually consistent with our discussion in Sec. 4.1. Second, we can also notice that integrating each module into the overall framework can respectively bring about improvement in ranking performance, which further verifies their effectiveness.